# Negative Sampling in Recommendation: A Survey and Future Directions

### v1

@@ 3.3. Adversarial Negative Generation
Adversarial Negative Generation (ANG) is a sophisticated technique which devise to enhance the robustness and performance in RS. Same as other negative sampling strategies, the core of ANG seeks to address the challenge of imbalanced training data, where positive interactions tend to dominate while real negative instances are relatively sparse. Existing ANG studies can be fundamentally categorized into two distinct paradigms: Generative ANG and Sampled ANG. The key emphases, advantages, and challenges of each strategy are systematically outlined in Table 4.
@@
Generative ANG endeavors to leverage a generative adversarial network (GAN) (Goodfellow et al., 2014), in which the generator and discriminator play a game-theoretical minimax game. The generator attempts to synthesize the plausible negative instances, while the discriminator seeks to distinguish between the genuine and generated samples. Through the iterative adversarial training, the generator becomes increasingly adept at generating mendacious but plausible negative instances, thereby enabling diverse negative generation, improving model generalization and substantially enriching the robustness of the recommender. However, it faces the risk of pattern breakdown and may lead to the relatively complex training process. On the other hand, Sampled ANG strategically samples appropriate items from the corpus with the Nash equilibrium state between the generator and discriminator in GAN. These selected items are then incorporated into the recommenders’ training process. Its advantage lies in guaranteeing the simplified GAN training procedure and the utilization of instances originating exclusively from authentic data. This eliminates any potential training bias stemming from varying levels of GAN’s generation complexity. Consequently, it effectively alleviates the imbalance between positive and negative samples in RS, enabling its training on a more balanced dataset. Nevertheless, it encounters challenges such as the homogenization of negatives and the difficulty in covering all potential user preferences. In the forthcoming sub-sections, we will expound upon these two methods from the existing ANG studies in RS.
3.3.1. Generative ANG
Generative ANG attempts to capture the underlying data distribution and generate the mendacious but plausible samples with an implicit generative model instead of imposing any assumption on the existing data. It can maintain the Nash equilibrium between generator and discriminator to generate negative samples that are similar to positive samples but have subtle differences, making them hard to distinguish from positive samples.(Wang et al., 2018b; Gao et al., 2019; Sun et al., 2020; Ding et al., 2019; Hao et al., 2021a; Xie et al., 2022b; Chen et al., 2022b; Zhou et al., 2021). Prevailing research in Generative ANS generally employs three methods for generating negative samples: user-specific generation, distribution-associated generation, and content-aware generation. Subsequently, we will elucidate each of these approaches in accordance with relevant works.
User-specific generation. Generative ANG leverages the adversarial mechanism to generate the targeted negative samples for different users by generators and maximize the objective function between real and generated item representations by discriminators, thereby pushing the recommender to its performance limit. RNS (Ding et al., 2019) designs an embedding-based sampler to generate exposure-alike negative instances and train a better recommender as the predictor through the feature matching technique and the adversarial training instead of directly choosing from exposure data. Different from RNS which only models the GAN as the sample generator, ELECRec (Chen et al., 2022b) proposes to define the recommendation as a discriminative task rather than a generative task, that is, trains the discriminator as the final recommender to distinguish if a sampled item is a ‘real’ target item or not. It defines the generator as an auxiliary model trained with the next-item prediction task to generative high-quality plausible samples, improving the discrimination ability of the discriminator from the intricate correlations.
Distribution-associated generation. Generative ANS also adopts the adversarial training paradigm to model the optimal data distribution from the correlation between users and items. For example, DRCGR (Gao et al., 2019) generates the relevancy distribution of targeted negative samples for different users with the generator and achieves a better estimate of the relevancy between the generated negative feedback sequence and the relevant positive feedback with the discriminator, thereby automatically learning the optimal recommendation strategies. To learn the true relevance distribution of users and items, PURE (Zhou et al., 2021) continuously samples on both users and items in the embedding space. Specifically, given a positive instance , the generator of PURE generates a fake item that is highly likely to be relevant to and a fake user that is likely to be relevant to from the random Gaussian noise.
3.3.2. Sampled ANG
Sampled ANG tries to fit the underlying distribution as much as possible, and samples the most likely items as negative instances to be connected with the positive sample (Wang et al., 2018b; Hao et al., 2021a; Wang et al., 2017b, 2018a). Different from the typical Generative ANG methods, the generator of Sampled ANG works directly select instances from the existing items in the corpus rather than their features. As one of the most famous Sampled ANG works, IRGAN (Wang et al., 2017b) stands as a pioneering endeavor that incorporates GANs (Goodfellow et al., 2014) into RS with a game-theoretical minimax game which optimizes both generator and discriminator iteratively. The primary objective of its generator lies in strategically selecting negative instances from the candidate pool and intermixing them with positive examples to confound the discriminator. Conversely, its discriminator is designed to distinguish positive samples from the negative instances sampled by the generator, and subsequently provide constructive feedback to refine the generator. Unlike the conventional Generative ANG methods which directly generate negative samples, IRGAN tries to sample negative instances based on the estimated correlation between candidate items and queries vie its generator. Similarly, the generator in GraphGAN (Wang et al., 2018a) attempts to approximate the underlying connectivity distribution in the graph structure and samples the relevant vertices similar to ground truth’s real immediate neighbors to deceive the discriminator. In contrast, its discriminator tries to detect whether the sampled vertex is from ground truth or selected by the generator. NMRN-GAN (Wang et al., 2018b) also develops a GAN-based noise sampler to adaptively sample the hardest negative examples from the randomly selected candidates, which considers both the specific user and the recommender parameters with the policy gradient-based reinforcement learning framework. In Multi-domain recommendation task, AFT (Hao et al., 2021a) proposes a domain-specific masked encoder and a GAN-based two-step feature translation to aggregate the domain-specific preferences from other domains, thereby retrieving the fake clicked samples in different domains for adversarial training.
@@ 3.4. Importance Re-weighting
Importance Re-weighting (IRW) is one of the sophisticated statistical techniques employed in Data Mining and Machine Learning (Chen et al., 2021a; Krishnan et al., 2020; Yu et al., 2020). Its core revolves around the nuanced adjustment of sample weights to prioritize the significance of certain negative samples within a given dataset (Chen et al., 2023d; Zhuo et al., 2022). It plays an integral role in scenarios where a dataset exhibits long-tailed problems or when specific samples possess varying degrees of importance. For the recommendation task where unobserved samples vastly outnumber positive feedback, IRW can be instrumental in rebalancing the real-world dataset by assigning diverse weights to different samples, thus enabling the algorithm to focus its learning efforts on the rare but crucial negative instances (Wang et al., 2021b; Chen et al., 2023b). The significance of IRW cannot be overstated. Its essential idea is consistent with the memorization effect: recommenders tend to learn accessible user preferences in the early stage of the training process and eventually memorize all complicated interactions. Properly assigning weights to individual samples attunes recommenders to the intricacies of imbalanced implicit datasets, thus recalibrating the learning process for the disparate data distribution. This adaptive weighting not only addresses the challenge of imbalanced data but also fosters the development of more robust and precise recommenders, ultimately bring in superior prediction accuracy.
The investigation of IRW algorithms classify them into Attention-based IRW, Knowledge-based IRW, and Debiased IRW. These strategies’ critical aspects, advantages, and challenges are systematically delineated in Table 5. Attention-based IRW emphasizes assigning diverse importance to each item with user interest attention, making it adaptably and capable of flexibly adjusting weights. However, it faces challenges due to its complexity and relatively poor interpretability. Knowledge-based IRW focuses on identifying each item’s importance with external structured knowledge, offering suitability for cold-start issues but introducing the additional dependency on knowledge. Debiased IRW aims to correct ubiquitous biases, assigning higher weights to overlooked items to provide equitable and diverse recommendations. Nevertheless, it involves the trade-off between fairness and precision. In the subsequent sub-sections, we will introduce the technical details of Attention-based IRW, Knowledge-based IRW, and Debiased IRW, respectively.
@@
| Attention-based IRW (Lian et al., 2020a; Wang et al., 2021c; Zhao et al., 2018; Gao et al., 2022a; Chen et al., 2021a; Covington et al., 2016; Wu et al., 2022a; Chen et al., 2023d; Zhuo et al., 2022) |
| Knowledge-based IRW (Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2021a; Wang et al., 2021b; Chen et al., 2019b) |
@@
3.4.1. Attention-based IRW
Attention-based IRW assigns importance scores and adjusts weights for each items, which could focus on the most relevant items for each user and provide suitable recommendations aligned with the user’s personalized interests with the attention mechanism. (Gao et al., 2022a; Chen et al., 2021a; Covington et al., 2016; Wu et al., 2022a; Chen et al., 2023d; Zhuo et al., 2022). To efficiently train recommenders with millions of items in the real-world application, this strategy attempt to sample negative candidates from the background distribution and then correct them via importance weighting. Firstly, a series of works attempt to focus on the safe (easy) negative samples while overlooking the hard negative samples. To adaptively prune the large-loss interactions along the training process, ADT (Wang et al., 2021c) devises the truncated loss paradigm to discard the hard samples (large-loss ones) with a dynamic threshold in each iteration and a reweighted loss paradigm to adaptively lower the weights of large-loss samples. UIB (Zhuo et al., 2022) innovatively introduces an auxiliary score for each user to define a personalized decision boundary and individually penalize samples that cross the boundary with a hybrid loss of the point-wise paradigm and the pair-wise paradigm. Moreover, some studies assume that more informative negative samples can contribute more to the gradient of loss function, the magnitude of the gradient becomes much larger so that training can be accelerated. To assist the learning process from the most important historical transitions, DEERS (Zhao et al., 2018) leverage the prioritized sampling strategy to greedily select the most informative negative sample while ensuring that the probability of being sampled is monotonic in a transition’s priority in reinforcement learning.
Recently, some related studies imitate the learning attention in human perception to assign different weights to diverse negative samples from simple to complex (whose informativeness is from less to more). PRIS (Lian et al., 2020a) proposes a personalized negative sampler from simple to complex based on rejection sampling to assign larger weights to more informative negative samples, which can contribute more to the gradient of the ranking loss function and accelerate the convergence rate. With the aim to address the distribution mismatch, REINFORCE (Chen et al., 2019a) takes the first-order approximation and ignores the state visitation differences as the importance weights of diverse negative samples with the Normalized Importance Sampling (NIS) strategy to construct a slightly biased estimator of the policy gradient with lower variance. SGDL (Gao et al., 2022a) first investigates the memorization effect within the representative recommenders’ learning process, delineating it into the noise-resistant period and the noise-sensitive period. Then, SGDL collects memorized interactions as denoising signals at the noise-resistant period and imposes a weight on each sample loss to enhance the robustness of the subsequent training during the noise-sensitive period. GeoSAN (Lian et al., 2020b) also proposes a weighted BCE loss based on importance sampling to reweight the unvisited locations with the negative probability, such that more informative locations can be more emphasized even with the uniform negative sampler. BIAO (Chen et al., 2023d) highlights the behavior-level effect in the source domain according to the corresponding item’s global impact and local user-specific impact during the proposed bi-level optimization process, which may have different importance for the cross-domain recommender’s optimization. ContraRec (Wang et al., 2023a) leverages the temperature hyper-parameter in the proposed CTC loss to assign diverse attention to negative samples according to their hardness in prediction.
3.4.2. Knowledge-based IRW
Knowledge-based IRW incorporates the external structured knowledge into the user’s preference modeling process to identify the importance of items from the corpus and assigns the corresponding weights in model optimization (Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2019b). The solutions employed in the Attention-based IRW studies design a weighting function that maps each sample’s score or loss value to its learning weight, thereby enabling the recommender to adaptively learn the information of each negative sample when model optimizing (Gao et al., 2022a; Chen et al., 2021a; Covington et al., 2016; Wu et al., 2022a). However, knowledge-based IRW argues that users’ social contexts and items’ heterogeneous information serve as influential determinants in shaping the reciprocal engagements between users and items (Yang et al., 2023b). In tandem, these elements covertly impact the confidence of the data. When an item is more popular among the user’s social neighbors, the user is more likely to know the item, and his feedback is more attributed to his real preference. Correspondingly, in instances where the items in the user’s historical behaviors exhibit a level of coherence in their content, recommender can effortlessly predict the items sharing the similar semantic attributes (Chen et al., 2021a; Wang et al., 2021b). This, in turn, addresses the user’s preference for particular content within the contextual paradigm, and this knowledge will be more reliable in deriving user preference.
Within the realm of recommendation, the pre-defined knowledge is typically categorized into user-oriented social knowledge and item-oriented heterogeneous knowledge. Different knowledge-based IRW methodologies are commonly utilized to integrate these distinct forms of knowledge into the process of importance re-weighting. For the user-oriented social knowledge, SamWalker (Chen et al., 2019b) simulates item information propagation along the real social network and adaptively specifies individual confidence weights to different interactions based on user’s social contexts. To mitigate SamWalker’s reliance on authentic user affiliations, SamWalker++ (Wang et al., 2021b) introduces the rationale of “wisdom of the crowds” to define the pseudo-social network, in which similar users are connected with specific item nodes or community nodes to improve the accuracy of learned data confidence. Under this scenario, SamWalker++ can model data confidence and assign adaptive confidence weights to each interaction without any side information. Furthermore, for the item-oriented heterogeneous knowledge, PREMERE (Kim et al., 2021a) proposes a meta-learning-based reweighting strategy to produce the most suitable weight at the current optimization state, which can handle the rarity of positive samples and the noisiness of negative samples simultaneously. URL (Chen et al., 2021a) leverages an auxiliary task to consider the user state and the context knowledge in reinforcement learning and reweights each example with the long-term reward to avoid the high variance in the learning of the policies caused by the inverse-propensity weighting. MMT-Net (Krishnan et al., 2020) propose an inverse novelty measure, noted as the context-bias, to attenuate the original supervised loss and decorrelate interactions by emulating variance-reduction, thereby inducing a novelty-weighted strategy focused on harder samples as training proceeds.
3.4.3. Debiased IRW
Debiased IRW identifies and corrects the ubiquitous biases that exist in RS (e.g., popularity bias, exposure bias) and assigns higher weights to items that have been overlooked in the past to deliver more equitable and diverse recommendations (Shen et al., 2021; Yu et al., 2020; Chen et al., 2023b). It operates on the premise that traditional recommenders may inadvertently exacerbate disparities by assigning undue weight to certain categories or user preferences.
By assigning diverse weights to items with various categories, debiased IRW endeavors to counteract any predispositions that may be inherent in the traditional recommenders. It is able to foster more equitable item representation, ensuring that no certain categories disproportionately influence the prediction results. For example, FairNeg (Chen et al., 2023b) proposes a negative sampling distribution mixup mechanism, which incorporates both the importance-aware negative sampling distribution and the fairness-aware negative sampling distribution to simultaneously strengthen the feature representations and item-oriented group fairness. Furthermore, some debiased IRW works also meticulously design and incorporate a suite of fairness metrics and concepts to evaluate the fairness and mitigate the inherent bias. SAR-Net (Shen et al., 2021) designs a concept of Fairness Coefficient (FC) to measure each sample’s importance and use it to reweigh the prediction in the proposed debiased expert networks, thereby mitigating the data fairness issue caused by manual intervention. IF4URec (Yu et al., 2020) introduces the Influence Function (IF) to measure the effect on the estimator when adding a small perturbation on one training sample, which could reveal the importance of one training sample. It is able to utilize the IF to reweight the training loss of each sample to get less loss in an unbiased validation for debiasing.

### v2

@@ 3.3. Adversarial Negative Generation
Adversarial Negative Generation (ANG) is a sophisticated technique which devise to enhance the robustness and performance of RS. Same as other negative sampling strategies, the core of ANG seeks to address the imbalanced data issue, where positive interactions tend to dominate while real negative instances are relatively sparse. Existing ANG studies can be fundamentally categorized into two distinct paradigms: Generative ANG and Sampled ANG. In the forthcoming sub-sections, we will expound upon these two methods from the existing ANG studies in RS.
3.3.1. Generative ANG
Generative ANG attempts to capture the underlying data distribution and generate the mendacious but plausible samples with a generative model instead of imposing any assumption on existing data. It can maintain the Nash equilibrium between generator and discriminator in generative adversarial network (GAN) (Goodfellow et al., 2014) to generate negative samples that are similar to positive samples but have subtle differences, making them hard to distinguish from positive samples (Wang et al., 2018b; Gao et al., 2019; Sun et al., 2020; Ding et al., 2019; Hao et al., 2021a; Xie et al., 2022b; Chen et al., 2022b; Zhou et al., 2021). Prevailing research in Generative ANS generally employs three methods for generating negative samples: user-specific generation, distribution-associated generation, and content-aware generation. Subsequently, we will elucidate each of them in accordance with relevant works.
User-specific generation. Generative ANG leverages the adversarial mechanism to generate the targeted negative samples for different users by generators and maximize the objective function between real and generated item representations by discriminators, thereby pushing the recommender to its performance limit. RNS (Ding et al., 2019) designs a reinforced negative sampler to generate exposure-alike negative instances and train a better recommender as the predictor through the feature matching technique and the adversarial training paradigm instead of directly choosing from exposure data. Different from RNS, ELECRec (Chen et al., 2022b) regards recommendation task as a discriminative task rather than a generative task, that is, training the discriminator as the final recommender to distinguish if a sampled item is a ‘real’ target item or not. It defines the generator as an auxiliary model of the next-item prediction task to generative high-quality plausible samples, improving the discrimination ability of the discriminator from the intricate correlations.
Distribution-associated generation. Generative ANS also adopts the adversarial training paradigm to model the optimal data distribution from the correlation between users and items. For example, DRCGR (Gao et al., 2019) generates the relevancy distribution of targeted negative samples for different users with the generator and achieves a better estimation of the relevancy between the generated negative feedback sequence and the relevant positive feedback with the discriminator, thereby automatically learning the optimal recommendation strategies. To learn the true relevance distribution of users and items, PURE (Zhou et al., 2021) continuously samples on both users and items in the representation space. Specifically, given a positive instance , the generator of PURE generates a fake item that is highly likely to be relevant to and a fake user that is likely to be relevant to from the random Gaussian noise.
3.3.2. Sampled ANG
Sampled ANG tries to fit the underlying distribution as much as possible, and sample the most likely items as negative instances from the captured distribution (Wang et al., 2018b; Hao et al., 2021a; Wang et al., 2017b, 2018a). Different from the typical Generative ANG methods, the generator of Sampled ANG works directly select instances from the existing items in the corpus rather than generating item representations. As one of the most famous Sampled ANG works, IRGAN (Wang et al., 2017b) stands as a pioneering endeavor that incorporates GANs into RS with a game theoretical minimax game which optimizes both generator and discriminator iteratively. The primary objective of its generator lies in strategically selecting negative instances from the candidate pool and intermixing them with positive examples to confound the discriminator. Conversely, its discriminator is designed to distinguish positive samples from the negative instances sampled by the generator, and subsequently provide constructive feedback to refine the generator. To be specific, IRGAN tries to sample negative instances based on the estimated correlation between candidate items and queries vie its generator. Similarly, the generator in GraphGAN (Wang et al., 2018a) attempts to approximate the underlying connectivity distribution in the graph structure and samples the relevant vertices similar to ground truth’s real immediate neighbors to deceive the discriminator. In contrast, its discriminator tries to detect whether the sampled vertex is from ground truth or selected by the generator. NMRN-GAN (Wang et al., 2018b) also develops a GAN-based noise sampler to adaptively sample the hardest negatives from the randomly selected candidates, which considers both the specific user and the recommender parameters with the policy gradient-based reinforcement learning framework. AFT (Hao et al., 2021a) proposes a domain-specific masked encoder and a GAN-based two-step feature translation to aggregate the domain-specific preferences from other domains, thereby retrieving the fake clicked samples in different domains for adversarial training.
@@ 3.3.2. Sampled ANG
Summary: We systematically outline the generation strategies, advantages, and challenges of each ANG strategy in Table 4. Here, Generative ANG endeavors to leverage GAN to generate the mendacious yet plausible samples, in which the generator and discriminator play a game-theoretical minimax game. The generator attempts to synthesize the plausible negative instances, while the discriminator seeks to distinguish between the genuine and generated samples. Through the iterative adversarial training, the generator becomes increasingly adept at generating mendacious but plausible negative instances, thereby enabling diverse negative generation, improving recommender generalization. However, it faces the risk of pattern breakdown and may lead to the relatively complex training process. In contrast, Sampled ANG strategically samples appropriate items from the corpus with the Nash equilibrium state between the generator and discriminator. Its advantage lies in guaranteeing the simplified GAN training procedure and the utilization of instances originating exclusively from authentic data. This eliminates any potential training bias stemming from varying levels of GAN’s generation complexity. Consequently, it effectively alleviates the imbalance between positive and negative samples in RS, enabling its training on a more balanced dataset. Nevertheless, it encounters challenges such as the homogenization of negatives and the difficulty in covering all potential user preferences.
@@ 3.4. Importance Re-weighting
Importance Re-weighting (IRW) is one of the sophisticated statistical techniques employed in Data Mining (Chen et al., 2021a; Krishnan et al., 2020; Yu et al., 2020; Wu and Ye, 2020). Its core revolves around the nuanced adjustment of sample weights to prioritize the significance of certain important negative samples (Chen et al., 2023d; Zhuo et al., 2022). For the recommendation task where unobserved samples vastly outnumber positive feedback, IRW can be instrumental in rebalancing the real-world dataset by assigning diverse weights to different samples, thus enabling the algorithm to focus its learning efforts on the rare but crucial negative instances (Wang et al., 2021b; Chen et al., 2023b). Its essential idea is consistent with the memorization effect: recommenders tend to learn accessible user preferences in the early stage of the training process and eventually memorize all complicated interactions. Properly assigning weights to individual samples attunes recommenders to the intricacies of imbalanced implicit datasets, thus recalibrating the learning process for the disparate data distribution. This adaptive weighting not only addresses the challenge of imbalanced data but also fosters the development of more robust and precise recommenders, ultimately bring in superior prediction accuracy. The investigation of IRW strategies classifies them into Attention-based IRW, Knowledge-aware IRW, and Debiased IRW and we will introduce their technical details in the subsequent sub-sections.
3.4.1. Attention-based IRW
Attention-based IRW assigns importance scores and adjusts weights for each item, which could focus on the most relevant items for each user and provide suitable recommendations aligned with the user’s personalized interests with the attention mechanism. (Gao et al., 2022; Chen et al., 2021a; Dai et al., 2024b; Chen et al., 2023d; Zhuo et al., 2022). To efficiently train recommenders with millions of items in the real-world applications, this strategy attempts to sample negative candidates from the uniform distribution and then correct them via importance weighting. A series of works attempt to focus on the safe (easy) negative samples while overlooking the hard negative samples. To adaptively prune the large-loss interactions, ADT (Wang et al., 2021c) devises the truncated loss paradigm to discard the large-loss samples (hard ones) with a dynamic threshold and a reweighted loss paradigm to adaptively lower the weights of these samples. UIB (Zhuo et al., 2022) innovatively introduces an auxiliary score for each user to define a personalized decision boundary and individually penalize samples that cross the boundary with a hybrid loss of point-wise and pair-wise paradigm. Moreover, some studies assume that more informative negative samples can contribute more to the gradient, the magnitude of the gradient becomes much larger so that training can be accelerated. To assist the learning process from the most important historical transitions, DEERS (Zhao et al., 2018) proposes a prioritized sampling strategy to greedily select the most informative negative sample while ensuring that the probability of being sampled is monotonic in a transition’s priority in reinforcement learning.
Recently, some studies imitate the learning attention in human perception to assign different weights to diverse negative samples from simple to complex (whose informativeness is from less to more). PRIS (Lian et al., 2020a) proposes a personalized negative sampler based on rejection sampling to assign larger weights to more informative negative samples, which can contribute more to the gradient of the ranking loss function and accelerate the convergence rate. To address the distribution mismatch, REINFORCE (Chen et al., 2019a) takes the first-order approximation and ignores the state visitation differences as the importance weights of diverse negative samples with the normalized importance sampling strategy to construct a slightly biased estimator of the policy gradient with lower variance. SGDL (Gao et al., 2022) investigates the memorization effect within several representative recommenders’ learning process, delineates them into the noise-resistant period and the noise-sensitive period. Then it collects memorized interactions as denoising signals at the noise-resistant period and imposes a weight on each sample loss to enhance the robustness of the noise-sensitive period. GeoSAN (Lian et al., 2020b) also proposes a weighted BCE loss based on importance sampling to reweight the unvisited locations with the negative probability, enabling more informative locations can be more emphasized even with the uniform negative sampler. ContraRec (Wang et al., 2023a) leverages the temperature hyper-parameter in the proposed CTC loss to assign diverse attention to negative samples according to their hardness in prediction.
3.4.2. Knowledge-aware IRW
Attention-based IRW studies typically design a reweighting function that maps each sample’s score or loss value to its learning weight, thereby enabling the recommender to adaptively learn the information of each negative sample in model optimization (Gao et al., 2022; Wu et al., 2022a). However, Knowledge-aware IRW argues that users’ social contexts and items’ heterogeneous information serve as influential determinants in shaping the reciprocal engagements between users and items (Yang et al., 2023b; Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2019b). In tandem, these elements covertly impact the confidence of the data. When an item is more popular among this user’s social neighbors, the user is more likely to know the item, and his feedback is more attributed to his real preference. Correspondingly, in the situation that user’s historical behaviors exhibit a level of coherence in their content, recommender can effortlessly predict the items sharing the similar semantic attributes (Chen et al., 2021a; Wang et al., 2021b). This, in turn, addresses the user’s preference for particular content within the contextual paradigm, and this knowledge will be more reliable in deriving user preference.
Within the realm of recommendation, the pre-defined knowledge is typically categorized into user-oriented social knowledge and item-oriented heterogeneous knowledge. Different knowledge-aware IRW methods are commonly utilized to integrate these distinct forms of knowledge into importance re-weighting. For the user-oriented social knowledge, SamWalker (Chen et al., 2019b) simulates item information propagation along the real social network and adaptively specifies individual confidence weights to different interactions based on user’s social contexts. To mitigate the reliance on authentic user affiliations, SamWalker++ (Wang et al., 2021b) introduces the rationale of “wisdom of the crowds” to define the pseudo-social network, in which similar users are connected with specific item nodes or community nodes. Under this scenario, SamWalker++ can model data confidence and assign adaptive confidence weight to each interaction without any side information. Furthermore, for the item-oriented heterogeneous knowledge, PREMERE (Kim et al., 2021a) proposes a meta-learning-based reweighting strategy to produce the most suitable weight at each optimization state, which can handle the rarity of positive samples and the noisiness of negative samples simultaneously. URL (Chen et al., 2021a) leverages an auxiliary task to consider user state and context knowledge in reinforcement learning and reweighs each item with a long-term reward to avoid the high variance in policy learning caused by the inverse-propensity weighting. MMT-Net (Krishnan et al., 2020) propose an inverse novelty measure, noted as context-bias, to attenuate the original supervised loss by emulating variance-reduction, thereby inducing a novelty-weighted strategy focused on harder samples as training proceeds.
3.4.3. Debiased IRW
Debiased IRW identifies and corrects the ubiquitous biases inherent in RS (e.g., popularity bias, exposure bias) and assigns higher weights to items that have been overlooked in the past to deliver more equitable and diverse recommendations (Shen et al., 2021; Yu et al., 2020; Chen et al., 2023b). It operates on the premise that traditional recommenders may inadvertently exacerbate disparities by assigning undue weight to certain categories or user preferences.
By assigning diverse weights to items with various categories, debiased IRW endeavors to counteract any predispositions within traditional recommenders. It can foster more equitable item representation, ensuring that no certain categories disproportionately influence the prediction results. For example, FairNeg (Chen et al., 2023b) proposes a negative sampling distribution mixup mechanism, which incorporates both importance-aware and fairness-aware sampling distribution to simultaneously strengthen the item representation and item-oriented group fairness. Furthermore, some debiased IRW works also meticulously design and incorporate a suite of fairness metrics and concepts to evaluate the fairness and mitigate the inherent bias. SAR-Net (Shen et al., 2021) designs a concept of “Fairness Coefficient” to measure each sample’s importance and use it to reweigh the prediction of the debiased expert networks, thereby mitigating the data fairness issue caused by manual intervention. IF4URec (Yu et al., 2020) introduces the “Influence Function” to measure the effect on the estimator when adding a small perturbation on training sample, which could reveal its importance. It is able to utilize “Influence Function” to reweigh the training loss of each sample to get less loss in an unbiased validation for debiasing.
@@ 3.4.3. Debiased IRW
| Attention-based IRW (Lian et al., 2020a; Wang et al., 2021c; Zhao et al., 2018; Gao et al., 2022; Chen et al., 2021a; Covington et al., 2016; Wu et al., 2022a; Chen et al., 2023d; Zhuo et al., 2022) |
| Knowledge-aware IRW (Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2021a; Wang et al., 2021b; Chen et al., 2019b) |
@@
Summary: We have systematically delineated the reweighting functions, advantages and challenges of these above IRW strategies in Table 5. Here, Attention-based IRW emphasizes assigning diverse importance to each item with user interest attention, making it adaptably and capable of flexibly adjusting weights. However, it faces challenges due to its complexity and relatively poor interpretability. Knowledge-aware IRW focuses on identifying each item’s importance with external structured knowledge, offering suitability for cold-start issues but introducing the additional dependency on knowledge. Debiased IRW aims to correct ubiquitous biases, assigning higher weights to overlooked items to provide equitable and diverse recommendations. Nevertheless, it involves the trade-off between fairness and precision.